# reference.bib

### v1

@@ 5. Operator $V_{N}$ and Dirichlet Series
We start by writing
$V_{N}\phi=N^{k/2-1}\sum_{A\in\Gamma_{1}\backslash M_{2}(\mathbb{Z})_{N}}\phi_{\sqrt{N}}|_{k,(\mathbb{Z}^{n},N\sigma)}\left(\frac{1}{\sqrt{N}}A\right),$
where we define $\phi_{\sqrt{N}}(\tau,z):=\phi\left(\tau,\sqrt{N}z\right)$. We can then follow the proof in [kohnen_skoruppa, pages 554-556] line by line and arrive at the result that the adjoint $V_{N}^{*}:J_{k}(\mathbb{Z}^{n},N\sigma)\longrightarrow J_{k}(\mathbb{Z}^{n},\sigma)$ is given by:
$\phi\longmapsto N^{k/2-2n-1}\sum_{X\mod{N\mathbb{Z}^{2n}}}\sum_{A\in\Gamma_{1}\backslash M_{2}(\mathbb{Z})_{N}}\phi_{\sqrt{N}^{-1}}|_{k,(\mathbb{Z}^{n},N\sigma)}\left(\frac{1}{\sqrt{N}}A\right)|_{k,(\mathbb{Z}^{n},N\sigma)}X.$
Let us now compute the action on the Fourier coefficients. We choose a set of representatives of the form $\begin{pmatrix}a&b\\ 0&d\end{pmatrix}$ with $ad=N$ and $0\leq b<d$. We then obtain that the above expression can be written as:
By then substituting the Fourier expansion:
for $\phi$, the above can be written as:
Now, the term
is zero unless $d\mid\frac{\frac{1}{2}qS^{-1}[r]-D}{qN}$ and $d\mid r$, meaning it divides all of its components. In that case, the sum equals $dN^{n}$. The conditions then imply that we can replace $r\longmapsto dr$ and $D\longmapsto Dd^{2}$. The last one follows from the fact that $N=ad$ and so $d\mid N$ as well. Hence, we obtain that the expression equals
$N^{k-n-1}\sum_{\lambda\in\mathbb{Z}^{n}/N\mathbb{Z}^{n}}\sum_{d\mid N}d^{1-k}\sum_{\begin{subarray}{c}D<0,r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r]\mod{\frac{qN}{d}}\end{subarray}}c_{\phi}(d^{2}D,dr)e\left(\left(\frac{\frac{1}{2}qS^{-1}[r]+\frac{1}{2}qS[\lambda]+qr^{t}\lambda-D}{q}\right)\tau+\right.\\ \left.+r^{t}z+\lambda^{t}Sz\right)=$
$=N^{k-n-1}\sum_{d\mid N}d^{1-k}\sum_{\lambda\in\mathbb{Z}^{n}/N\mathbb{Z}^{n}}\sum_{\begin{subarray}{c}D<0,r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r-S\lambda]\mod\frac{qN}{d}\end{subarray}}c_{\phi}(d^{2}D,d(r-S\lambda))e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{q}\tau+r^{t}z\right),$
after setting $r\longmapsto r-S\lambda$. Now, as in [kohnen_skoruppa], we set
$\lambda\equiv t+\frac{N}{d}t^{\prime}\mod N,$
with $t\mod\frac{N}{d}$ and $t^{\prime}\mod d$. We then have
$S^{-1}(d(r-S\lambda))\equiv S^{-1}(d(r-St))\mod N\mathbb{Z}^{n}$
and
$D\equiv\frac{1}{2}qS^{-1}[r-St]\mod\frac{qN}{d}.$
The first property can be seen easily and the second follows from the fact that we already have $D\equiv\frac{1}{2}qS^{-1}[r-S\lambda]\mod\frac{qN}{d}$ and also we can check:
So, it suffices to show that
$\frac{1}{2}S[t]\equiv\frac{1}{2}S[\lambda]\mod{\frac{N}{d}}.$
But we can write $\lambda=t+\frac{N}{d}\alpha$ with $\alpha\in\mathbb{Z}^{n}$ and we can then see the claim follows.
Hence, because of Remark 1, the final expression we obtain has the form (after replacing $d$ by $N/d$):
$\sum_{d\mid N}d^{k-(n+1)}\sum_{t\in\mathbb{Z}^{n}/d\mathbb{Z}^{n}}\sum_{\begin{subarray}{c}D<0,r\in\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[r-St]\mod qd\end{subarray}}c_{\phi}\left(\left(\frac{N}{d}\right)^{2}D,\frac{N}{d}(r-St)\right)e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{q}\tau+r^{t}z\right).$
What remains to prove now (after fixing $D,r$ with the appropriate conditions) is
$=\sum_{d\mid N}d^{k-(n+1)}\sum_{\begin{subarray}{c}t\mod dS\mathbb{Z}^{n}\\ D\equiv\frac{1}{2}qS^{-1}[t]\mod{qd}\end{subarray}}c_{\phi}\left(\left(\frac{N}{d}\right)^{2}D,\frac{N}{d}t\right).$
This follows by setting $u=r-St$ and then observing that $D\equiv\frac{1}{2}qS^{-1}[u]\mod qd$,
and by using the fact that $c_{\phi}(D,s)=c_{\phi}(D,s^{\prime})$ if $s\equiv s^{\prime}\mod NS\mathbb{Z}^{n}$. So we can consider the entries $r-St$ $\mod NS\mathbb{Z}^{n}$ and all of these are different $\mod dS\mathbb{Z}^{n}$. ∎
[…]
$\mathcal{P}_{k,D,r}(\tau^{\prime},z,\tau):=\sum_{N\geq 1}V_{N}P_{k,D,r}(\tau,z)e(N\tau^{\prime}),$
which is then a cusp form of weight $k$ with respect to $\Gamma_{S}$ (see [jacobi_lattice, Theorem 2.3 (i)]) and an element of the Maass space (cf. [sugano_lifting, Lemma 6.1] and Remark 3). Here, by abusing notation, we use $P_{k,D,r}$ to actually denote the Poincaré series $\frac{1}{\lambda_{k,D}}P_{k,D,r}$, with the quantities defined in Section 4. We then define the Dirichlet series:
@@
A corollary of [jacobi_lattice, Theorem 2.3 (i)] is that $S_{k}(L,\sigma)$ is generated by $\{P_{k,D,r}\mid(D,r)\in\widetilde{\textup{supp}}(L,\sigma)\}$ (cf. [jacobi_lattice, Corollary 2.4]). This observation, together with Theorem 3.1, gives us that the Maass space $M_{k}^{*}(\Gamma_{S})$, is generated by $\{\mathcal{P}_{k,D,r}\mid(D,r)\in\widetilde{\textup{supp}}(L,\sigma)\}$.
Let us now relate the coefficients $c_{\phi_{N}}\left(\frac{N^{2}}{d^{2}}D,\frac{N}{d}s\right)$ with the Fourier coefficients of $F$. $F$ admits a Fourier expansion of the form
$F(Z)=\sum_{\tilde{r}\in L_{0}^{*}}A(\tilde{r})e(r^{t}S_{0}Z)=\sum_{N=1}^{\infty}\phi_{N}(\tau,z)e(N\tau^{\prime}),$
where $Z=(\tau^{\prime},z,\tau)$ and $\tilde{r}=(n,r,N)$ with $r\in L^{*}$. We can then write
$\phi_{N}(\tau,z)=\sum_{n\in\mathbb{Z},r\in L^{*}}A(n,r,N)e(n\tau-r^{t}Sz)=\sum_{n\in\mathbb{Z},r\in\mathbb{Z}^{n}}A(n,S^{-1}r,N)e(n\tau-r^{t}z).$
But now $\begin{pmatrix}1_{2}&&\\ &-1_{n}&\\ &&1_{2}\end{pmatrix}\in\Gamma_{S}$ and therefore if $Z=(\tau^{\prime},z,\tau)\in\mathcal{H}_{S}$, we have
$F\left((\tau^{\prime},-z,\tau)\right)=F((\tau^{\prime},z,\tau)),$
which then implies $A(n,r,N)=A(n,-r,N)$ for all $n,N\in\mathbb{Z},r\in L^{*}$. Therefore, after setting $r\longmapsto-r$, we can re-write the above as:
$\phi_{N}(\tau,z)=\sum_{n\in\mathbb{Z},r\in\mathbb{Z}^{n}}A(n,S^{-1}r,N)e(n\tau+r^{t}z).$
Comparing now the above expression with the expression
$\sum_{D,r}c_{\phi_{N}}\left(D,r\right)e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{qN}\tau+r^{t}z\right),$
we see that we after setting $r\longmapsto\frac{N}{d}s,D\longmapsto\frac{N^{2}}{d^{2}}D$, we obtain
$c_{\phi_{N}}\left(\frac{N^{2}}{d^{2}}D,\frac{N}{d}s\right)=A\left(\frac{N}{d}\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)\right).$
From the above considerations, we arrive at the following proposition:
@@
The proof follows from Proposition 5.1 and Proposition 4.1. ∎
@@ 6. Relation to the class number
In this section, our goal is to bring $\mathcal{D}_{F,\mathcal{P}_{k,D,r}}$ into a form similar to the one in [kohnen_skoruppa, page 553]. We first need some definitions of the adelized groups and of genus and classes of lattices.
@@
$V_{\mathbb{A}}=\left\{v\in\prod_{p\leq\infty}V_{p}\mid v_{p}\in L_{p}\textup{ for almost all }p<\infty\right\}$
$\textup{GL}(V)_{\mathbb{A}}=\left\{\alpha\in\prod_{p\leq\infty}\textup{GL}(V_{p})\mid\alpha_{p}L_{p}=L_{p}\textup{ for almost all }p<\infty\right\}$
@@
Let $G\subset GL_{n}(\mathbb{Q})$ any algebraic linear group. For any field $K$ containing $\mathbb{Q}$, we denote by $G_{K}$ the group of $K-$rational points in $G$. Now, for any prime number $p$, we abbreviate by $G_{p}$ the group $G_{\mathbb{Q}_{p}}$. Let also $G_{\infty}$ denote $G_{\mathbb{R}}$. We then define
@@
Let $L$ denote a $\mathbb{Z}-$lattice in $V$ and $G\subset GL_{n}(\mathbb{Q})$ any algebraic group. Then, for any $x\in G$, $xL$ is also a $\mathbb{Z}-$lattice in $V$. We define the class of $L$ to be $\{xL\mid x\in G\}$. Similarly, for any $x\in G_{\mathbb{A}}$, $xL$ is again a $\mathbb{Z}-$lattice in $V$. The set $\{xL\mid x\in G_{\mathbb{A}}\}$ is called the genus of $L$.
Now, the genus of $L$ can be decomposed into disjoint union of classes. If $C=\{x\in G_{\mathbb{A}}\mid xL=L\}$, the map $xC\longmapsto xL$ gives a bijection between $G_{\mathbb{A}}/C$ and the genus of $L$, so gives a bijection between $G\backslash G_{\mathbb{A}}/C$ and the set of classes contained in the genus of $L$. In general, if $U$ is an open subgroup of $G_{\mathbb{A}}$, we call the number #$\#(G\backslash G_{\mathbb{A}}/U)$ the class number of $G$ with respect to $U$.
Now, with all the main definitions out of the way, we want to deal with the vectors appearing in Proposition 5.2. We start with the next lemma:
@@
for all $d\geq 1$ coprime to $D$. Here, for a $\mathbb{Z}-$lattice $\Lambda$ in $V_{0}$, $\mathfrak{b}$ a fractional ideal of $\mathbb{Q}$ and $q\in\mathbb{Q}^{\times}$, we define:
$\Lambda[q,\mathfrak{b}]=\{x\in V_{0}\mid\phi_{0}[x]=q\textup{ and }\phi_{0}(x,\Lambda)=\mathfrak{b}\}.$
@@
Firstly, for any vector $\xi\in\Xi_{d}$, we can directly compute $\phi_{0}[\xi]=-\frac{D}{q}$. It now remains to show that the $\mathbb{Z}-$ideal
$\{\xi^{t}S_{0}x\mid x\in\mathbb{Z}^{n+2}\}$
equals $\mathbb{Z}$ (because the bilinear form is $\phi_{0}=\frac{1}{2}S_{0}$). But for any basis vector $e_{i}$ of the lattice $\mathbb{Z}^{n+2}$, we have
$\xi^{t}S_{0}e_{i}=\begin{cases}d&\textup{if }i=1\\ -s_{i-1}&\textup{if }2\leq i\leq n+1\\ \frac{\frac{1}{2}qS^{-1}[s]-D}{qd}&\textup{if }i=n+2\end{cases}$
so the above ideal is contained in $\mathbb{Z}$ by the conditions that define the set $\Xi_{d}$. Now, if it were equal to $k\mathbb{Z}$ for some $k\geq 1$, then we would have $k\mid d$, $k\mid s$ and $k\mid\frac{\frac{1}{2}qS^{-1}[s]-D}{qd}$ which would then imply $k\mid D$, and so $k=1$ as $d,D$ are assumed to be coprime. ∎
We now have the following very important lemma:
@@
The class number of $G_{\mathbb{Q}}^{*}$, defined as #$\#(G^{*}_{\mathbb{Q}}\backslash G^{*}_{\mathbb{A}}/C)$, where
$C=\{x\in G^{*}_{\mathbb{A}}\mid xL_{0}=L_{0}\},$
@@
The proof follows from [shimura2004arithmetic, Theorem 9.26], as the ideal class group of $\mathbb{Q}$ is trivial and $G^{*}_{\mathbb{R}}$ is not compact because we are in the indefinite case (the signature is $(1,n+1)$). Hence, $Y$ in that theorem is the whole $\mathbb{Q}_{\mathbb{A}}^{\times}$ and so $J_{0}$, hence $J_{\phi}$, is the whole $J$, in the notation of the theorem. Also $n\geq 3$ there corresponds to $n\geq 1$ in our notation. ∎
Fix now an element $\xi\in L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$ and consider the algebraic subgroup of $G_{\mathbb{Q}}^{*}$:
@@
We note that $H(\xi)_{\mathbb{Q}}=SO^{\psi}(W)$, where $W:=\{x\in V_{0}\mid\phi_{0}(x,\xi)=0\}$ and $\psi:=\phi_{0}\mid_{W}$.
We now have the following Proposition, which is a special case of [shimura_diophantine_2006, Theorem 2.2]:
@@
$L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})\longleftrightarrow H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C),$
@@
If $k\in L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$, then there is some $\alpha\in G^{*}_{\mathbb{Q}}$ such that $k=\alpha\xi$ (Witt’s theorem, [shimura2004arithmetic, Lemma 1.5 (ii)]). We then assign the $H(\xi)_{\mathbb{Q}}-$class of $\alpha^{-1}L_{0}$ to $k$. In particular, this then gives
#$\#\left(L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})\right)=\#(H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)).$
@@
The remaining assertions can be seen from [shimura_diophantine_2006, (2.7)]. In the notation there, $D$ is the set $C$ of Lemma 6.2 and therefore we can take $Y$ to be the trivial set in [shimura_diophantine_2006, Theorem 2.2, (iii)]. This then gives exactly the result claimed, because $G_{\mathbb{Q}}^{*}\cap D=G_{\mathbb{Q}}^{*}\cap C=\Gamma(L_{0})$ of Section 2. ∎
@@
$L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]=\bigsqcup_{i=1}^{h}\Gamma(L_{0})\xi_{i}.$

### v2

@@ 6. Dirichlet Series and relation with Fourier coefficients
The proof is similar to [13, Lemma 1]. We will show for $N\geq 1$ that
$\left\langle\phi_{N},\psi_{N}\right\rangle=\mathcal{O}(N^{k}),$
with the constant depending only on $F,G$. Indeed, fix $(\tau,z)\in\mathbb{H}\times\mathbb{C}^{n}$ and write $\tau=u+iv,\textup{ }z=x+iy$. If $F(q)=\sum_{N=1}^{\infty}\phi_{N}(\tau,z)q^{N}$, with $q=e^{2\pi i\tau^{\prime}}$, we have by Cauchy’s integral formula that
$\phi_{N}(\tau,z)=\oint_{|q|=r}\frac{F(q)}{q^{N+1}}\textup{d}q,$
for any $0<r<e^{-\pi S[y]/v}$. The bounds follow from the fact that $S_{0}[\textup{Im}Z]>0$ (here $Z=(\tau^{\prime},z,\tau)\in\mathcal{H}_{S}$). If now $\tau^{\prime}=u^{\prime}+iv^{\prime}$, the integral can be written as
$\phi_{N}(\tau,z)=\int_{0}^{1}F(Z)e^{-2\pi iN\tau^{\prime}}\textup{d}u^{\prime},$
for any $v^{\prime}>S[y]/2v$. But now $\left|F(Z)\right|\left(S_{0}[\textup{Im}Z]/2\right)^{k/2}$ is bounded on $\mathcal{H}_{S}$ from [11, II, Lemma 3.28], say by a constant $C>0$. Therefore, after choosing $v^{\prime}=S[y]/2v+1/N$, we have
$\left|\phi_{N}(\tau,z)\right|\leq Ce^{2\pi}\int_{0}^{1}\left(S_{0}[\textup{Im}Z]/2\right)^{-k/2}e^{\pi NS[y]/v}\textup{d}u^{\prime}=Ce^{2\pi}\left(\frac{v}{N}\right)^{-k/2}e^{\pi NS[y]v}.$
Similarly for $\psi_{N}$ and then the claim follows from the definition of the inner product in Definition 3.5. ∎
From now on, we let $k>n/2+2$ be an even integer. Let then $F\in S_{k}(\Gamma_{S})$ and write $\phi_{N}$ for its Fourier–Jacobi coefficients. For $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$, let also
$\mathcal{P}_{k,D,r}(\tau^{\prime},z,\tau):=\sum_{N\geq 1}V_{N}P_{k,D,r}(\tau,z)e(N\tau^{\prime}).$ (8)
From [28, Corollary 6.7], we have that $\mathcal{P}_{k,D,r}\in S_{k}^{*}(\widetilde{\Gamma}_{S})$ (see Definition 2.6). Here, by abusing notation, we write $P_{k,D,r}$ to actually denote the Poincaré series $\displaystyle{\frac{1}{\lambda_{k,D}}P_{k,D,r}}$, with the quantities defined in Proposition 5.1. Then, we have the Dirichlet series
@@
Even though $F$ and $\mathcal{P}_{k,D,r}$ are taken invariant with respect to different modular groups, we note here that the proof of Lemma 6.1 is still valid, because $F\in S_{k}(\Gamma_{S})\subset S_{k}(\widetilde{\Gamma}_{S})$.
Remark 5.
A corollary of [17, Theorem 2.3 (i)] is that $S_{k}(L,\sigma)$ is generated by $\left\{P_{k,D,r}\mid(D,r)\in\widetilde{\textup{supp}}(L,\sigma),\ S^{-1}r\pmod{L}\right\}$ (cf. [17, Corollary 2.4]). This observation, together with [28, Corollary 6.7] gives us that the Maass space $S_{k}^{*}(\widetilde{\Gamma}_{S})$ is generated by $\left\{\mathcal{P}_{k,D,r}\mid(D,r)\in\widetilde{\textup{supp}}(L,\sigma),\ S^{-1}r\pmod{L}\right\}$. It is therefore enough to consider $G$ to be a Poincaré series in (7).
We will now relate the $N$th term of $\mathcal{D}_{F,\mathcal{P}_{k,D,r}}(s)$ with the Fourier coefficients of $F$.
@@
$F$ admits a Fourier expansion of the form (see equation (2))
$F(Z)=\sum_{\tilde{r}\in L_{0}^{*}}A(\tilde{r})e({\widetilde{r}}^{t}S_{0}Z)=\sum_{N=1}^{\infty}\phi_{N}(\tau,z)e(N\tau^{\prime}),$
where $Z=(\tau^{\prime},z,\tau)\in\mathcal{H}_{S}$ and $\tilde{r}=(m,r,N)$ with $r\in L^{*}$. We can then write
$\phi_{N}(\tau,z)=\sum_{m\in\mathbb{Z},r\in L^{*}}A(m,r,N)e(m\tau-r^{t}Sz)=\sum_{m\in\mathbb{Z},r\in\mathbb{Z}^{n}}A(m,S^{-1}r,N)e(m\tau-r^{t}z).$
But now $\begin{pmatrix}-1_{2}&&\\ &1_{n}&\\ &&-1_{2}\end{pmatrix}\in\Gamma_{S}$ and therefore if $Z=(\tau^{\prime},z,\tau)\in\mathcal{H}_{S}$, we have
$F\left((\tau^{\prime},-z,\tau)\right)=(-1)^{k}F((\tau^{\prime},z,\tau))=F((\tau^{\prime},z,\tau)),$
which then implies $A(m,r,N)=A(m,-r,N)$ for all $m,N\in\mathbb{Z},\ r\in L^{*}$. Therefore, after setting $r\longmapsto-r$, we can re-write the above as
$\phi_{N}(\tau,z)=\sum_{m\in\mathbb{Z},r\in\mathbb{Z}^{n}}A(m,S^{-1}r,N)e(m\tau+r^{t}z).$ (9)
Now, a priori, we can write
$\phi_{N}(\tau,z)=\sum_{D,r}c_{\phi_{N}}\left(D,r\right)e\left(\frac{\frac{1}{2}qS^{-1}[r]-D}{qN}\tau+r^{t}z\right),$ (10)
with $(D,r)$ as in equation (4). Hence, from Proposition 4.1 and Proposition 5.1, we have
$\displaystyle\langle V_{N}^{*}\phi_{N},P_{k,D,r}\rangle$ $\displaystyle=\sum_{d\mid N}d^{k-(n+1)}\sum_{\begin{subarray}{c}s\ (\mathrm{mod}\ dS\mathbb{Z}^{n})\\ D\equiv\frac{1}{2}qS^{-1}[s]\ (\mathrm{mod}\ qd)\end{subarray}}c_{\phi}\left(\left(\frac{N}{d}\right)^{2}D,\frac{N}{d}s\right)$
$\displaystyle=\sum_{d\mid N}d^{k-(n+1)}\sum_{\begin{subarray}{c}s\ (\mathrm{mod}\ dS\mathbb{Z}^{n})\\ D\equiv\frac{1}{2}qS^{-1}[s]\ (\mathrm{mod}\ qd)\end{subarray}}A\left(\frac{N}{d}\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)\right),$
because after setting $r\longmapsto Ns/d,D\longmapsto N^{2}D/d^{2}$ in (10), we obtain from (9)
$c_{\phi_{N}}\left(\frac{N^{2}}{d^{2}}D,\frac{N}{d}s\right)=A\left(\frac{N}{d}\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)\right).\qed$
@@ 7. Relation to the class number
In this Section, our goal is to bring $\mathcal{D}_{F,\mathcal{P}_{k,D,r}}(s)$ into a form similar to the one in [13, page 553]. We first need some definitions of the adelized groups and of genus and classes of lattices.
@@
$V_{\mathbb{A}}=\left\{v\in\prod_{p\leq\infty}V_{p}\mid v_{p}\in L_{p}\textup{ for almost all }p<\infty\right\},$
$\textup{GL}(V)_{\mathbb{A}}=\left\{\alpha\in\prod_{p\leq\infty}\textup{GL}(V_{p})\mid\alpha_{p}L_{p}=L_{p}\textup{ for almost all }p<\infty\right\}.$
@@
Let $G\subset\hbox{GL}_{n}(\mathbb{Q})$ any algebraic linear group. For any field $K$ containing $\mathbb{Q}$, we denote by $G_{K}$ the group of $K$-rational points in $G$. For any prime number $p$, we abbreviate by $G_{p}$ the group $G_{\mathbb{Q}_{p}}$. Let also $G_{\infty}$ denote $G_{\mathbb{R}}$. We then define
@@
Let $L$ denote a $\mathbb{Z}$-lattice in $V$ and $G\subset\hbox{GL}_{n}(\mathbb{Q})$ any algebraic group. Then, for any $x\in G$, $xL$ is also a $\mathbb{Z}$-lattice in $V$. We define the class of $L$ to be $\{xL\mid x\in G\}$. Similarly, for any $x\in G_{\mathbb{A}}$, $xL$ is again a $\mathbb{Z}$-lattice in $V$. The set $\{xL\mid x\in G_{\mathbb{A}}\}$ is called the genus of $L$.
[…]
Now, with all the main definitions out of the way, we want to deal with the vectors appearing in Proposition 6.1. For a $\mathbb{Z}$-lattice $\Lambda$ in $V_{0}$, $\mathfrak{b}$ a fractional ideal of $\mathbb{Q}$ and $q\in\mathbb{Q}^{\times}$, we define
$\Lambda[q,\mathfrak{b}]:=\left\{x\in V_{0}\mid\phi_{0}[x]=q\textup{ and }\phi_{0}(x,\Lambda)=\mathfrak{b}\right\}.$
We then have the following Lemma.
@@
for all $d\geq 1$ coprime to $D$.
@@
Firstly, for any vector $\xi\in\Xi_{d}$, we can directly compute $\phi_{0}[\xi]=-D/q$. It now remains to show that the $\mathbb{Z}$-ideal
$\left\{\xi^{t}S_{0}x\mid x\in\mathbb{Z}^{n+2}\right\}$
equals $\mathbb{Z}$ (because the bilinear form is $\phi_{0}=S_{0}/2$). But for any basis vector $e_{i}$ of the lattice $\mathbb{Z}^{n+2}$, we have
$\xi^{t}S_{0}e_{i}=\begin{cases}d&\textup{if }i=1\\ -s_{i-1}&\textup{if }2\leq i\leq n+1\\ \frac{\frac{1}{2}qS^{-1}[s]-D}{qd}&\textup{if }i=n+2\end{cases},$
so the above ideal is contained in $\mathbb{Z}$ by the conditions that define the set $\Xi_{d}$. Now, if it were equal to $k\mathbb{Z}$ for some $k\geq 1$, then we would have
$k\mid d,\ k\mid s,\textup{ and }k\mid\frac{\frac{1}{2}qS^{-1}[s]-D}{qd}.$
This would then imply $k\mid D$, and so $k=1$, as $d,\ D$ are assumed to be coprime. ∎
We now have the following very important Lemma.
@@
The class number of $G_{\mathbb{Q}}^{*}$, defined as $\#(G^{*}_{\mathbb{Q}}\backslash G^{*}_{\mathbb{A}}/C)$, where
$C=\{x\in G^{*}_{\mathbb{A}}\mid xL_{0}=L_{0}\},$ (11)
@@
This is shown in [26, Remark 2.4, (5)], which is an improvement of [24, Theorem 9.26], as the technical assumptions are weakened. ∎
Assume now it exists $\xi\in L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$ and consider the algebraic subgroup of $G_{\mathbb{Q}}^{*}$
@@
We note that if $W:=\{x\in V_{0}\mid\phi_{0}(x,\xi)=0\}$ and $\psi:=\phi_{0}\mid_{W}$, then
$H(\xi)_{\mathbb{Q}}=\hbox{SO}^{\psi}(W)=\{\alpha\in\textup{GL}(W)\mid\psi[\alpha w]=\psi[w],\ \forall w\in W\}.$
We now have the following Proposition, which is a special case of [26, Theorem 2.2].
@@
[…]
@@
If $k\in L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$, then there is some $\alpha\in G^{*}_{\mathbb{Q}}$ such that $k=\alpha\xi$ (Witt’s theorem, [24, Lemma 1.5 (ii)]). We then assign the $H(\xi)_{\mathbb{Q}}$-class of $\alpha^{-1}L_{0}$ to $k$. In particular, this then gives
[…]
@@
The remaining assertions follow from [26, Theorem 2.2, (iv)], because $G_{\mathbb{Q}}^{*}$ has class number one (Lemma 7.2) and $G_{\mathbb{Q}}^{*}\cap C=\Gamma(L_{0})$. ∎
@@
$L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]=\bigsqcup_{i=1}^{h}\Gamma(L_{0})\xi_{i}.$ (12)